A worker will lay out wood frames in two rectangles, and will pour concrete into these frames for a small patio. The length of each rectangle is 2 feet more than its width. if the total length of the wood to be used is 64 feet, find the LEGNTH and WIDTH of the rectangles. ( ignore the thickness of the wood)
Answer by mananth(16946)

A worker will lay out wood frames in two rectangles, and will pour concrete into these frames for a small patio. The length of each rectangle is 2 feet more than its width. if the total length of the wood to be used is 64 feet, find the LEGNTH and WIDTH of the rectangles. ( ignore the thickness of the wood)

Let the width be x
Length = x+3
2*(4x+4)=64
8x+8=64
8x=56
X=7
Length =9